# Analyzing Probabilistic Logic Shields for Multi-Agent Reinforcement Learning

### v2

@@ B.4.2 Experimental conditions
The agents observe the grid as a concatenated (flattened) vector consisting of the set of one-hot encoded vectors for each position on the grid to describe what is in contained in them: either ($i$) nothing, ($ii$) a stag, ($iii$) a plant, ($iv$) the observing agent, ($v$) the other agent, or ($vi$) both agents. Thus, the full input is a vector of size is $5\times 5\times 6=150$.
@@ C.3 Extended Public Goods Game
Let us instead create a shield that that uses a model $\hat{\mu}$ of the true $\mu$ estimated over training so far. Additionally, we can use the standard deviation of the multiplication factors so far, $\hat{\sigma}$ to model the uncertainty the agent has with respect to the current $f_{t}$ compared to $\hat{\mu}$. This can be done online without the need of a buffer using Welford’s method [37]. We then compare $\hat{\mu}$ to the predetermined Nash outcomes and decide if it is high enough to warrant defecting – if $\hat{\mu}\geq 1$, cooperate, else defect. In Shield 5, this is set as the (Boolean) value of the fact sensor(mu_high) – its value is set to 1 if $\hat{\mu}\geq 1$ else 0. The agents will err on the side of a cooperative strategy when $21<\hat{\mu}<$ 2, where the rational strategy would be mixed.
We then might want to include uncertainty to increase exploitation if it has high certainty about the environment (align the agent more strongly with the inductive bias induced via the shield), and contrariwise, increase exploration if there is high uncertainty (guide the agent more weakly). We must thus construct a smooth distance measure $d$ between $f_{t}$ and $\hat{\mu}$ bounded between $[0,1]$ such that $d(f_{t},\hat{\mu})=1$ when $f_{t}=\hat{\mu}$ and $d(f_{t},\hat{\mu})=0$ when $|f_{t}-\hat{\mu}|\rightarrow\infty$. One of many possibilities is described here. First, we compute the $z$-score of the current $f_{t}$ with respect to the estimated parameters so far:
@@
We then use the CDF of a standard normal distribution to translate this into a probability value $\Phi(z(f_{t}\mid\hat{\mu},\hat{\sigma}))$. Subtracting 0.5 centers $\Phi(z(f_{t}\mid\hat{\mu},\hat{\sigma}))$ around 0 for $z(f_{t}\mid\hat{\mu},\hat{\sigma})=0$. Doubling the absolute value and subtracting from 1 inverts the scale to make it 1 at the mean and symmetrically 0 at extreme tails. Thus the final distance function is:
@@
In Shield 5 below, the value of the fact sensor(f_certainty) is thus set to $d(f_{t},\hat{\mu})$. Thus using the inputs defined above, the shield is constructed as follows:
@@ Appendix D Single-Agent Experiment: CartSafe
CartSafe is modification of OpenAI Gym’s classical control problem CartPole [6] with simple constraints inspired by jemaw [18]. The problem has the following elements: the environment consists of a rod attached to a box or cart at one end and is allowed to freely swing around the attachment point. This setup exists in a bounded 2-dimensional real space with a set width and an active physics simulation. At each time step, the agent has two actions: accelerate the cart to the left or to the right. The agent observes the cart’s position, velocity, the angular velocity of the rod and the angle of the rod in radians. The goal is to actively balance the rod upright without leaving the bounds of the environment. The agent receives a reward at each time step based on the angle of the pole ($\theta_{pole}$) with respect to vertical: $r^{t}=1$ if $-15^{\circ}\leq\theta_{pole}\leq 15^{\circ}$ and $r^{t}=-1$ otherwise. The environment terminates when a maximum number of steps ($t_{max}$) are reached or the cart moves off screen. The additional constraints that are added to this setup are such that the cart must not enter the region bounded by the leftmost and rightmost quarters of the environment – thus the agent must learn to balance the rod near the center of the screen.
@@ D.2 Shield construction
[…]
@@ Appendix F Training Curves
The training curves for Stag-Hunt (Figure 13), 2-player Extended Public Goods Game (Figure 14) 2-player Markov Stag-Hunt (Figures 15 and 16) are given below. All remaining experiment training curves will be available via a Weights and Biases dashboard for convenience at the publication of this paper, or will be provided at the request of the reader.
Figure 13: Training, evaluation and safety results for Stag-Hunt for PLPG-based agents. Results are smoothed using a rolling average with window size 50. The lines represent the mean and the shadow the standard deviation over 5 seeds.

### v3

@@ B.4.2 Experimental conditions
The agents observe the grid as a concatenated (flattened) vector consisting of the set of one-hot encoded vectors for each position on the grid to describe what is in contained in them: either ($i$) nothing, ($ii$) a stag, (iiiiii) a plant, ($iv$) the observing agent, ($v$) the other agent, or ($vi$) both agents. Thus, the full input is a vector of size is $5\times 5\times 6=150$.
@@ C.3 Extended Public Goods Game
Let us instead create a shield that that uses a model $\hat{\mu}$ of the true $\mu$ estimated over training so far. Additionally, we can use the standard deviation of the multiplication factors so far, $\hat{\sigma}$ to model the uncertainty the agent has with respect to the current $f_{t}$ compared to $\hat{\mu}$. This can be done online without the need of a buffer using Welford’s method [44]. We then compare $\hat{\mu}$ to the predetermined Nash outcomes and decide if it is high enough to warrant defecting – if $\hat{\mu}\geq 1$, cooperate, else defect. In Shield 5, this is set as the (Boolean) value of the fact sensor(mu_high) – its value is set to 1 if $\hat{\mu}\geq 1$ else 0. The agents will err on the side of a cooperative strategy when $1<\hat{\mu}<2$, where the rational strategy would be mixed.
[…]
@@
We then use the CDF of a standard normal distribution to translate this into a probability value $)\Phi(z(f_{t}\mid\hat{\mu},\hat{\sigma}$)). Subtracting 0.5 centers $)\Phi(z(f_{t}\mid\hat{\mu},\hat{\sigma}$)) around 0 for $z(f_{t}\mid\hat{\mu},\hat{\sigma})=0$. Doubling the absolute value and subtracting from 1 inverts the scale to make it 1 at the mean and symmetrically 0 at extreme tails. Thus the final distance function is:
@@
[…]
@@ Appendix D Single-Agent Experiment: CartSafe
CartSafe is modification of OpenAI Gym’s classical control problem CartPole [7] with simple constraints inspired by jemaw [22]. The problem has the following elements: the environment consists of a rod attached to a box or cart at one end and is allowed to freely swing around the attachment point. This setup exists in a bounded 2-dimensional real space with a set width and an active physics simulation. At each time step, the agent has two actions: accelerate the cart to the left or to the right. The agent observes the cart’s position, velocity, the angular velocity of the rod and the angle of the rod in radians. The goal is to actively balance the rod upright without leaving the bounds of the environment. The agent receives a reward at each time step based on the angle of the pole ($\theta_{pole}$) with respect to vertical: $r^{t}=1$ if $15^{\circ}\leq\theta_{pole}\leq 15^{\circ}$ and $r^{t}=-1$ otherwise. The environment terminates when a maximum number of steps ($t_{max}$) are reached or the cart moves off screen. The additional constraints that are added to this setup are such that the cart must not enter the region bounded by the leftmost and rightmost quarters of the environment – thus the agent must learn to balance the rod near the center of the screen.
@@ D.2 Shield construction
This means that if the cart is within the permissible central region of the environment, the safety probability of the next action is computed as $(\texttt{sensor(cost)}=0)\implies(\texttt{safe\_next}=1)$, and the policy remains unchanged. If the cart is too far right, outside the permissible region, it is unsafe to move the cart towards the right. The right action gets progressively more unsafe as the cart gets further to the right. The safe policy will proportionally favor the left action as a result. Analogous behavior occurs if the cart is too far left instead of right.
@@ Appendix F Training Curves
The training curves for 2-player Extended Public Goods Game (Figure 13) and 2-player Markov Stag-Hunt (Figures 14 and 15) are given below. All remaining experiment training curves will be provided at the request of the reader.